OpenStudy (anonymous):

8 • [5 + (18 – 7)] ÷ 4

OpenStudy (bunnyzrockgirl):

32 i think

OpenStudy (anonymous):

WAIT

OpenStudy (bunnyzrockgirl):

WHAT? lol

OpenStudy (anonymous):

Which property justifies this statement? If x + y = z and z = x + 4, then x + y = x + 4. A. distributive B. associative C. symmetric D. transitive

OpenStudy (bunnyzrockgirl):

i dont know about that one. sorry!